### v2

@@ 4.6 Vehicle-to-Everything (V2X) Datasets
[luo2025mixed] Mixed-Signals 2025 Australia 0 45.1K LiDAR, BBox No Yes No Web
[xiang2025v2x] V2X-ReaLO 2025 China 25,028 25,028 RGB, LiDAR, ROS-bags Yes Yes No N/A
[wang2024deepaccident] DeepAccident 2024 Carla 57K 57K RGB, LiDAR, BEW Yes Yes No Git
[karvat2024adver] Adver-City 2024 Carla 24,087 24,087 RGB, LiDAR, IMU/GNSS Yes Yes No Web
[li2024multi] Multi-V2X 2024 Carla 549K 146K RGB, LiDAR, BBox Yes Yes No Git
[chen2024whales] WHALES 2024 Carla 70K 17K RGB, LiDAR, Trajectory Yes Yes No Git
[huang2024v2x] V2X-R 2024 China 150,908 37,727 RGB, LiDAR, BBox Yes Yes Yes Git
[zhou2024v2xpnp] V2XPnP 2024 USA 208K 40K RGB, BBox, Trajectory Yes Yes Yes Web
[gamerdinger2024scope] SCOPE 2024 Carla 17,600 17,600 RGB, LiDAR, BBox Yes Yes No Git
[wang2024rcdn] OPV2V-N 2024 Carla 11,464 11,464 RGB, LiDAR, BBox Yes Yes No N/A
[li2022v2x] V2X-Sim 2022 Carla 10K 10K RGB, LiDAR, GPS, Depth Yes Yes Yes Git
[xu2022v2x] V2X-ViT 2022 Carla 11,447 11,447 RGB, LiDAR, BBox Yes Yes No Git
[mao2022dolphins] DOLPHINS 2022 Carla 42,376 42,376 RGB, LiDAR Yes Yes No Web
[xiang2024v2x] V2X-Real 2022 USA 171K 34K RGB, LiDAR, BBox Yes Yes No Web
@@ 4.6.1 Simulation-based Datasets
DeepAccident [wang2024deepaccident], Adver-City [karvat2024adver], Multi-V2X [li2024multi], WHALES [chen2024whales], SCOPE [gamerdinger2024scope], OPV2V-N [wang2024rcdn], V2X-Sim [li2022v2x], V2X-ViT [xu2022v2x], and DOLPHINS [mao2022dolphins] datasets are created using simulation platforms, offering precise ground-truth labels, synchronized multi-agent data, and diverse cooperative scenarios. DeepAccident dataset [wang2024deepaccident], released in 2024 by HKU, Huawei, and Dalian University of Technology, is a safety-focused synthetic dataset for accident prediction in V2X environments. It was generated by CARLA and features 57,000 frames and 285,000 samples of multi-agent interactions in 12 major intersection accident types. Each scene includes data from RGB cameras and LiDAR sensors on four vehicles and one roadside unit. Tasks include 3D detection, tracking, BEV segmentation, motion forecasting, and full-scene accident prediction. While highly structured and valuable for risk-aware learning, it lacks real-world sensor noise and requires heavy computation.
@@
The Adver-City dataset [karvat2024adver], released in March 2025 by Queen’s University, is a large-scale synthetic dataset for evaluating cooperative perception in adverse weather. Built using CARLA and OpenCDA, it includes 24,000 frames and 890,000 annotations across 110 scenarios featuring rain, fog, glare, and varying lighting and traffic conditions. Inspired by U.S. National Highway Traffic Safety Administration (NHTSA) crash data, it focuses on dangerous driving areas like intersections and rural curves. Data comes from five agents (ego, RSUs, and vehicles) equipped with RGB, segmentation cameras, LiDAR, and GNSS/IMU. It supports tasks like detection, tracking, and segmentation, with a label format matching OPV2V. Strengths include diverse agents, weather types (including glare), and realistic layouts. Limitations include a lack of real-world sensor data. Multi-V2X [li2024multi] is a large-scale synthetic cooperative perception benchmark released in 2024 by Tsinghua University. It is the first dataset specifically designed to assess autonomous driving performance under varying CAV (Connected and Autonomous Vehicle) penetration rates. Built using CARLA and SUMO, it includes 549,000 RGB images, 146,000 LiDAR frames, and 4.2 million annotated 3D bounding boxes across six object classes. The dataset simulates both V2V and V2I scenarios across urban, suburban, and rural settings. Vehicles and RSUs are equipped with RGB cameras, LiDARs, and GNSS modules. The six included towns feature diverse road topologies and traffic conditions. Multi-V2X supports 3D object detection, tracking, and benchmarking using fusion strategies (early, late, intermediate). Meanwhile, its synthetic nature limits realism and weather variation.
@@
WHALES (Wireless enHanced Autonomous vehicles with Large number of Engaged agentS) [chen2024whales] is a large-scale synthetic dataset released in 2024 by Tsinghua University. It is generated using CARLA, and features 70,000 RGB images, 17,000 LiDAR frames, and over 2 million annotated 3D bounding boxes, with an average of 8.4 agents per scene. WHALES supports tasks like object detection and CP benchmarking under realistic communication constraints. While it offers rich multimodal annotations and novel scheduling tasks, its synthetic nature limits realism, and its high computational demands may restrict accessibility. SCOPE (Synthetic COllective PErception) [gamerdinger2024scope] is a 2024 synthetic multi-agent dataset from the University of Tübingen and FZI Karlsruhe, designed to advance collective perception in AVs. It includes over 17,600 frames across 40+ scenarios with up to 24 agents, featuring realistic sensor models (RGB, solid-state LiDAR), detailed weather simulation, and urban environments. SCOPE offers annotations for object detection and semantic segmentation, supporting research on perception resilience, sensor alignment, and domain adaptation. Its strengths include high-fidelity environmental modeling and diverse road user types. Limitations include simulation-based constraints on real-world unpredictability and the high computational cost of processing multimodal data.
@@
V2X-Sim [li2022v2x] is a large-scale synthetic dataset released in 2022 by NYU, USC, and Shanghai Jiao Tong University. Built using CARLA and SUMO, it simulates V2X scenarios with synchronized multi-agent sensor data from LiDAR, RGB, semantic, and depth cameras on both vehicles and infrastructure. V2X-Sim’s strengths lie in its early contribution to multimodal, multi-agent CP benchmarks and its compatibility with open-source tools. However, its simulated nature limits realism in terms of sensor noise and unpredictable driving behavior. DOLPHINS (Dataset for cOLlaborative Perception enabled Harmonious and INterconnected Self-driving) [mao2022dolphins] is a large synthetic benchmark released in 2022 by Tsinghua University to support both V2V and V2I collaborative perception. It features synchronized full-HD images and 64-line LiDAR point clouds collected from ego vehicles, auxiliary vehicles, and RSUs across six diverse scenarios, including urban, highway, mountain, and rural settings under weather conditions like rain and fog. While DOLPHINS excels in scope and detail, it is limited by its synthetic nature, low frame rate for high-speed scenes, and lack of radar data.
@@ 4.6.2 Real-World Datasets
Mixed-Signals [luo2025mixed], V2X-ReaLO [xiang2025v2x], V2X-R [huang2024v2x], V2XPnP [zhou2024v2xpnp], and V2X-Real [xiang2024v2x] datasets are collected from physical vehicle deployments in real-world settings. Mixed-Signals [luo2025mixed] is a real-world V2X dataset released in 2025 by Cornell and the University of Sydney. It features data from three vehicles and one RSU with different LiDAR setups, as well as 45,100 point clouds and 240,600 3D bounding boxes. Its main strengths are showcasing sensor heterogeneity and real-world multi-agent perception. Limitations include a single location, short recording duration, and lack of camera or radar data. V2X-ReaLO [xiang2025v2x] is a real-world dataset released in March 2025 by UCLA. Built on V2X-Real, it offers 25,028 frames (6,850 annotated keyframes) from two connected vehicles and two RSUs, supporting real-time evaluation of early, late, and intermediate fusion strategies. Its key strength lies in practical and real-time V2X fusion under real-world conditions, with open-source support for modular experimentation. Limitations include a lack of radar data and limited geographic diversity.
V2X-R [huang2024v2x] is a multimodal synthetic dataset released in March 2025 to advance robust autonomous driving in adverse weather. It includes 12,079 driving scenarios with 37,727 LiDAR and radar point clouds, 150,908 RGB images, and over 170,000 3D bounding boxes. Additionally, it enables enhanced 3D object detection under fog and snow. Despite being synthetic and resource-intensive, V2X-R offers a critical platform for developing weather-resilient, multi-agent, multimodal autonomous perception systems. V2XPnP [zhou2024v2xpnp] is a large-scale real-world benchmark introduced by University of California researchers in December 2024. It includes around 40,000 frames across 100 vehicle-centric and 63 infrastructure-centric scenarios. Its advantages is combining temporal consistency with full V2X coverage, which is ideal for studying real-time multi-agent collaboration. However, it lacks detailed sensor specifications and demands high computational resources.
@@ 4.7 Infrastructure-to-Infrastructure (I2I) Datasets
The RCOOPER dataset [hao2024rcooper] was collected from simulated urban scenarios using CARLA and was released in 2020. The sensor setup included LiDAR, a camera on the vehicle, and LiDAR on the roadside infrastructure. The authors used the data to evaluate a cooperative perception system for 3D object detection using early and late fusion. It is one of the premier studies on cooperative perception. However, using synthetic data limits real-world applicability. There is no radar in the sensor setup. The InScope dataset [zhang2024inscope], released in 2024, was developed to address occlusion challenges in autonomous driving. It includes 187,787 annotated 3D bounding boxes and 303 tracking trajectories. The dataset supports four benchmark tasks: collaborative 3D object detection, multi-source fusion, domain adaptation, and multi-object tracking. Its ability to mitigate occlusion through infrastructure-side sensing is one of its main advantages; however, its drawbacks include limited geographic diversity and access requirements.
@@
[hao2024rcooper] RCooper 2024 China 50K 30K RGB, LiDAR, BBox Yes Yes No Git
[zhang2024inscope] InScope 2024 China 0 21,317 LiDAR, BBox, Trajectory No Yes No Git
@@ 5 Object Detection in Autonomous Vehicles
In 2D camera-based detection, algorithms rely on RGB imagery to identify and localize objects using bounding boxes projected onto the image plane [boroujeni2024ic]. These methods can be further classified by their underlying architecture into three main groups: CNN-based (e.g., YOLOv4, Faster R-CNN), Transformer-based (e.g., DETR, Deformable DETR), and Hybrid models that integrate convolutional backbones with transformer modules (e.g., Swin Transformer + Faster R-CNN). At the same time, 2D detection approaches can also be categorized by their input configuration, including monocular (single-view RGB), depth-based monocular, multi-view geometry-based, multi-camera system-level, and stereo camera setups, each tailored to specific perception requirements and environmental conditions. In contrast, 3D LiDAR-based detection methods operate on raw or preprocessed point clouds to generate precise spatial representations of objects. These techniques encompass different detection strategies, including anchor-based one-stage detectors (e.g., SECOND, YOLO3D), anchor-based two-stage detectors (e.g., PV-RCNN, Voxel R-CNN), and anchor-free detectors (e.g., 3DSSD, YOLOv8-3D), each offering trade-offs between speed, localization accuracy, and robustness to sparse data. In parallel, 3D detection approaches can also be categorized by their input setup, including point-based, voxel-based, pillar-based, BEV-based, point-voxel-based, and range-based. 2D–3D fusion detection combines visual and geometric information from both cameras and LiDAR sensors to enhance object recognition under challenging conditions. Fusion strategies are typically classified as early-fusion, mid-fusion, late-fusion, or special-fusion (not strictly fitting into the previous three). Additionally, they can be categorized based on their input modalities to three different categories. Lastly, LLM/VLM-based detection approaches utilize large language and vision-language models to incorporate textual context into visual reasoning, enabling semantic-level understanding and generalization across unseen categories. These techniques include zero-shot detection, grounded detection (region-text alignment), caption-supervised Detection (language-augmented supervision), multimodal pretrained detection (Joint vision-language models), prompt-based fine-tuning (domain-adapted queries), and scene-level understanding. This categorization provides a unique framework and valuable insight into the current landscape of object detection in AVs, guiding future research toward more robust and generalizable perception systems [elrod2025graph].
@@
[liang2022aspectnet] AspectNet 2022 ResNet-50 Mono 96.16 93.87 88.17 87.17 76.57 68.42 86.15 77.11 71.24
@@
[choi2019gaussian] Gaussian YOLOv3 2019 Darknet-53 Mono 90.61 90.20 81.19 87.84 79.57 72.30 89.31 81.30 80.20
@@
[choi2019gaussian] Gaussian YOLOv3 2019 Darknet-53 Mono 98.74 90.48 89.47 87.85 79.96 76.81 90.08 86.59 81.09
@@
[yi2019assd] ASSD 2019 ResNet-101 Mono 89.28 89.95 82.11 69.07 62.49 60.18 75.23 76.16 72.83
@@
[zhang2018single] RefineDet 2018 VGG-16 Mono 98.96 90.44 88.82 84.40 77.44 73.52 86.33 80.22 79.15
@@
[wu2017squeezedet] SqueezeDet 2017 SqueezeNet Mono 90.20 84.70 73.90 77.10 68.30 65.80 82.90 75.40 72.10
@@
From Table 14, we provide an in-depth discussion of two selected models: AspectNet [liang2022aspectnet] and SINet [hu2018sinet], which achieved the highest performance among all evaluated methods, and Gaussian YOLOv3 [choi2019gaussian], one of the most widely used models in the context of autonomous driving. AspectNet [liang2022aspectnet] approaches object detection tasks uniquely with an anchor-free detection scheme. It sets itself apart from comparable models by adding a special head that predicts object aspect ratios. Rather than tackline classification and regression features separately, the network combines them before reaching the aspect module, which allows geometric details to play a role in both areas. This additional branch is designed with just a few convolutional layers, ensuring it does not slow inference time too much. Another notable change is how positive and negative training samples are defined: the criteria for matching are tweaked to tackle the class imbalance that can hurt performance in single-stage models. These architectural decisions enhance the detector’s ability to recognize elongated or skewed objects, as the receptive field is shaped better to fit the actual geometry of the detected objects.
@@
Gaussian YOLOv3 [choi2019gaussian] keeps the overall layout of YOLOv3 intact, with three detection scales, feature aggregation from the backbone, and dense predictions. However, it changes the way bounding boxes are parameterized. Each coordinate is modeled as a Gaussian variable with its variance, meaning the network outputs not only position estimates but also an uncertainty term for each. These variances are fed into a modified regression loss based on negative log-likelihood, and they also influence the suppression of overlapping boxes at inference. The result is a system that can flag low-confidence localizations without discarding high-confidence detections that fall near ambiguous boundaries. Compared with AspectNet’s emphasis on encoding geometric priors into a deterministic pipeline, Gaussian YOLOv3 takes a probabilistic route to preserve the speed of a one-stage detector while offering a measure of confidence that can be applied to downstream tasks.
@@
Table 15 focuses on 3D object detection methods that use camera-based KITTI data, but generate 3D bounding boxes from 2D image inputs. From Table 15, we provide a detailed discussion of four selected models: M3D-RPN [brazil2019m3d], recognized as one of the most widely adopted models in this field; MonoDGP [Pu_2025_CVPR], which is the latest released method; MonoTAKD [Liu_2025_CVPR], which achieved the highest results in $AP_{BEV}$ CarCar and $Pedestrian$ evaluations; and MonoDiff [Ranasinghe_2024_CVPR] which achieved the highest results in $AP_{3D}$ CarCar evaluation on the KITTI dataset. It should be mentioned that the values highlighted in red represent the best performance achieved for each evaluation metric across all listed methods.
@@
The Monocular 3D Region Proposal Network (M3D-RPN) [brazil2019m3d] presents an integrated approach for generating 2D and 3D object proposals from a single RGB image. Built upon a DenseNet-121 backbone modified with dilated convolutions to preserve a stride of 16, the network branches into two complementary pathways: one employing standard spatially invariant convolutions to capture global context, and another utilizing depth-aware convolutions to encode row-dependent geometric cues. In the latter, feature maps are divided into horizontal bins, with each bin processed by distinct convolutional kernels to reflect perspective variations inherent to urban driving scenes. Anchors are designed to carry 2D and 3D parameters, initialized from dataset-derived statistics to provide strong geometric priors. Predictions from the two branches are combined to regress full 3D bounding box parameters, followed by a lightweight post-optimization step that refines object orientation through projection consistency between 3D boxes and their 2D counterparts. M3D-RPN reduces complexity while maintaining accuracy by consolidating proposal generation into a single end-to-end stage.
@@
MonoDGP [Pu_2025_CVPR] employs a transformer-based architecture that unites geometric priors with a novel perspective-invariant error formulation for depth refinement. A ResNet-50 backbone extracts multi-scale features, which are enhanced by a Region Segmentation Head (RSH) to emphasize foreground content and embed segment labels distinguishing object regions from background. This network’s detection is decoupled into two parallel decoding processes: a 2D visual decoder for initializing queries and reference points from appearance cues and a 3D depth-guided decoder that integrates visual and depth features to model spatial relationships. Rather than relying on direct depth prediction or multi-branch fusion, MonoDGP computes geometric depth using the projection formula and introduces an explicit geometry error term to correct systematic offsets caused by perspective and object surface effects. With its stable and compact statistical distribution, this error term reduces training complexity while improving depth accuracy. Combining query decoupling allows the method to achieve improved convergence stability and precise 3D localization without requiring auxiliary data sources.

### v3

@@ 4.6 Vehicle-to-Everything (V2X) Datasets
[158] Mixed-Signals 2025 Australia 0 45.1K LiDAR, BBox No Yes No Web
[262] V2X-ReaLO 2025 China 25,028 25,028 RGB, LiDAR, ROS-bags Yes Yes No N/A
[242] DeepAccident 2024 Carla 57K 57K RGB, LiDAR, BEW Yes Yes No Git
[111] Adver-City 2024 Carla 24,087 24,087 RGB, LiDAR, IMU/GNSS Yes Yes No Web
[123] Multi-V2X 2024 Carla 549K 146K RGB, LiDAR, BBox Yes Yes No Git
[37] WHALES 2024 Carla 70K 17K RGB, LiDAR, Trajectory Yes Yes No Git
[98] V2X-R 2024 China 150,908 37,727 RGB, LiDAR, BBox Yes Yes Yes Git
[319] V2XPnP 2024 USA 208K 40K RGB, BBox, Trajectory Yes Yes Yes Web
[72] SCOPE 2024 Carla 17,600 17,600 RGB, LiDAR, BBox Yes Yes No Git
[241] OPV2V-N 2024 Carla 11,464 11,464 RGB, LiDAR, BBox Yes Yes No N/A
[131] V2X-Sim 2022 Carla 10K 10K RGB, LiDAR, GPS, Depth Yes Yes Yes Git
[270] V2X-ViT 2022 Carla 11,447 11,447 RGB, LiDAR, BBox Yes Yes No Git
[165] DOLPHINS 2022 Carla 42,376 42,376 RGB, LiDAR Yes Yes No Web
[261] V2X-Real 2022 USA 171K 34K RGB, LiDAR, BBox Yes Yes No Web
@@ 4.6.1 Simulation-based Datasets
DeepAccident [242], Adver-City [111], Multi-V2X [123], WHALES [37], SCOPE [72], OPV2V-N [241], V2X-Sim [131], V2X-ViT [270], and DOLPHINS [165] datasets are created using simulation platforms, offering precise ground-truth labels, synchronized multi-agent data, and diverse cooperative scenarios. DeepAccident dataset [242], released in 2024 by HKU, Huawei, and Dalian University of Technology, is a safety-focused synthetic dataset for accident prediction in V2X environments. It was generated by CARLA and features 57,000 frames and 285,000 samples of multi-agent interactions in 12 major intersection accident types. Each scene includes data from RGB cameras and LiDAR sensors on four vehicles and one roadside unit. Tasks include 3D detection, tracking, BEV segmentation, motion forecasting, and full-scene accident prediction. While highly structured and valuable for risk-aware learning, it lacks real-world sensor noise and requires heavy computation.
@@
The Adver-City dataset [111], released in March 2025 by Queen’s University, is a large-scale synthetic dataset for evaluating cooperative perception in adverse weather. Built using CARLA and OpenCDA, it includes 24,000 frames and 890,000 annotations across 110 scenarios featuring rain, fog, glare, and varying lighting and traffic conditions. Inspired by U.S. National Highway Traffic Safety Administration (NHTSA) crash data, it focuses on dangerous driving areas like intersections and rural curves. Data comes from five agents (ego, RSUs, and vehicles) equipped with RGB, segmentation cameras, LiDAR, and GNSS/IMU. It supports tasks like detection, tracking, and segmentation, with a label format matching OPV2V. Strengths include diverse agents, weather types (including glare), and realistic layouts. Limitations include a lack of real-world sensor data. Multi-V2X [123] is a large-scale synthetic cooperative perception benchmark released in 2024 by Tsinghua University. It is the first dataset specifically designed to assess autonomous driving performance under varying CAV (Connected and Autonomous Vehicle) penetration rates. Built using CARLA and SUMO, it includes 549,000 RGB images, 146,000 LiDAR frames, and 4.2 million annotated 3D bounding boxes across six object classes. The dataset simulates both V2V and V2I scenarios across urban, suburban, and rural settings. Vehicles and RSUs are equipped with RGB cameras, LiDARs, and GNSS modules. The six included towns feature diverse road topologies and traffic conditions. Multi-V2X supports 3D object detection, tracking, and benchmarking using fusion strategies (early, late, intermediate). Meanwhile, its synthetic nature limits realism and weather variation.
@@
WHALES (Wireless enHanced Autonomous vehicles with Large number of Engaged agentS) [37] is a large-scale synthetic dataset released in 2024 by Tsinghua University. It is generated using CARLA, and features 70,000 RGB images, 17,000 LiDAR frames, and over 2 million annotated 3D bounding boxes, with an average of 8.4 agents per scene. WHALES supports tasks like object detection and CP benchmarking under realistic communication constraints. While it offers rich multimodal annotations and novel scheduling tasks, its synthetic nature limits realism, and its high computational demands may restrict accessibility. SCOPE (Synthetic COllective PErception) [72] is a 2024 synthetic multi-agent dataset from the University of Tübingen and FZI Karlsruhe, designed to advance collective perception in AVs. It includes over 17,600 frames across 40+ scenarios with up to 24 agents, featuring realistic sensor models (RGB, solid-state LiDAR), detailed weather simulation, and urban environments. SCOPE offers annotations for object detection and semantic segmentation, supporting research on perception resilience, sensor alignment, and domain adaptation. Its strengths include high-fidelity environmental modeling and diverse road user types. Limitations include simulation-based constraints on real-world unpredictability and the high computational cost of processing multimodal data.
@@
V2X-Sim [131] is a large-scale synthetic dataset released in 2022 by NYU, USC, and Shanghai Jiao Tong University. Built using CARLA and SUMO, it simulates V2X scenarios with synchronized multi-agent sensor data from LiDAR, RGB, semantic, and depth cameras on both vehicles and infrastructure. V2X-Sim’s strengths lie in its early contribution to multimodal, multi-agent CP benchmarks and its compatibility with open-source tools. However, its simulated nature limits realism in terms of sensor noise and unpredictable driving behavior. DOLPHINS (Dataset for cOLlaborative Perception enabled Harmonious and INterconnected Self-driving) [165] is a large synthetic benchmark released in 2022 by Tsinghua University to support both V2V and V2I collaborative perception. It features synchronized full-HD images and 64-line LiDAR point clouds collected from ego vehicles, auxiliary vehicles, and RSUs across six diverse scenarios, including urban, highway, mountain, and rural settings under weather conditions like rain and fog. While DOLPHINS excels in scope and detail, it is limited by its synthetic nature, low frame rate for high-speed scenes, and lack of radar data.
@@ 4.6.2 Real-World Datasets
Mixed-Signals [158], V2X-ReaLO [262], V2X-R [98], V2XPnP [319], and V2X-Real [261] datasets are collected from physical vehicle deployments in real-world settings. Mixed-Signals [158] is a real-world V2X dataset released in 2025 by Cornell and the University of Sydney. It features data from three vehicles and one RSU with different LiDAR setups, as well as 45,100 point clouds and 240,600 3D bounding boxes. Its main strengths are showcasing sensor heterogeneity and real-world multi-agent perception. Limitations include a single location, short recording duration, and lack of camera or radar data. V2X-ReaLO [262] is a real-world dataset released in March 2025 by UCLA. Built on V2X-Real, it offers 25,028 frames (6,850 annotated keyframes) from two connected vehicles and two RSUs, supporting real-time evaluation of early, late, and intermediate fusion strategies. Its key strength lies in practical and real-time V2X fusion under real-world conditions, with open-source support for modular experimentation. Limitations include a lack of radar data and limited geographic diversity.
V2X-R [98] is a multimodal synthetic dataset released in March 2025 to advance robust autonomous driving in adverse weather. It includes 12,079 driving scenarios with 37,727 LiDAR and radar point clouds, 150,908 RGB images, and over 170,000 3D bounding boxes. Additionally, it enables enhanced 3D object detection under fog and snow. Despite being synthetic and resource-intensive, V2X-R offers a critical platform for developing weather-resilient, multi-agent, multimodal autonomous perception systems. V2XPnP [319] is a large-scale real-world benchmark introduced by University of California researchers in December 2024. It includes around 40,000 frames across 100 vehicle-centric and 63 infrastructure-centric scenarios. Its advantages is combining temporal consistency with full V2X coverage, which is ideal for studying real-time multi-agent collaboration. However, it lacks detailed sensor specifications and demands high computational resources.
@@ 4.7 Infrastructure-to-Infrastructure (I2I) Datasets
The RCOOPER dataset [80] was collected from simulated urban scenarios using CARLA and was released in 2020. The sensor setup included LiDAR, a camera on the vehicle, and LiDAR on the roadside infrastructure. The authors used the data to evaluate a cooperative perception system for 3D object detection using early and late fusion. It is one of the premier studies on cooperative perception. However, using synthetic data limits real-world applicability. There is no radar in the sensor setup. The InScope dataset [308], released in 2024, was developed to address occlusion challenges in autonomous driving. It includes 187,787 annotated 3D bounding boxes and 303 tracking trajectories. The dataset supports four benchmark tasks: collaborative 3D object detection, multi-source fusion, domain adaptation, and multi-object tracking. Its ability to mitigate occlusion through infrastructure-side sensing is one of its main advantages; however, its drawbacks include limited geographic diversity and access requirements.
@@
[80] RCooper 2024 China 50K 30K RGB, LiDAR, BBox Yes Yes No Git
[308] InScope 2024 China 0 21,317 LiDAR, BBox, Trajectory No Yes No Git
@@ 5 Object Detection in Autonomous Vehicles
In 2D camera-based detection, algorithms rely on RGB imagery to identify and localize objects using bounding boxes projected onto the image plane [21]. These methods can be further classified by their underlying architecture into three main groups: CNN-based (e.g., YOLOv4, Faster R-CNN), Transformer-based (e.g., DETR, Deformable DETR), and Hybrid models that integrate convolutional backbones with transformer modules (e.g., Swin Transformer + Faster R-CNN). At the same time, 2D detection approaches can also be categorized by their input configuration, including monocular (single-view RGB), depth-based monocular, multi-view geometry-based, multi-camera system-level, and stereo camera setups, each tailored to specific perception requirements and environmental conditions. In contrast, 3D LiDAR-based detection methods operate on raw or preprocessed point clouds to generate precise spatial representations of objects. These techniques encompass different detection strategies, including anchor-based one-stage detectors (e.g., SECOND, YOLO3D), anchor-based two-stage detectors (e.g., PV-RCNN, Voxel R-CNN), and anchor-free detectors (e.g., 3DSSD, YOLOv8-3D), each offering trade-offs between speed, localization accuracy, and robustness to sparse data. In parallel, 3D detection approaches can also be categorized by their input setup, including point-based, voxel-based, pillar-based, BEV-based, point-voxel-based, and range-based. 2D–3D fusion detection combines visual and geometric information from both cameras and LiDAR sensors to enhance object recognition under challenging conditions. Fusion strategies are typically classified as early-fusion, mid-fusion, late-fusion, or special-fusion (not strictly fitting into the previous three). Additionally, they can be categorized based on their input modalities to three different categories. Lastly, LLM/VLM-based detection approaches utilize large language and vision-language models to incorporate textual context into visual reasoning, enabling semantic-level understanding and generalization across unseen categories. These techniques include zero-shot detection, grounded detection (region-text alignment), caption-supervised Detection (language-augmented supervision), multimodal pretrained detection (Joint vision-language models), prompt-based fine-tuning (domain-adapted queries), and scene-level understanding. This categorization provides a unique framework and valuable insight into the current landscape of object detection in AVs, guiding future research toward more robust and generalizable perception systems [66].
@@
[140] AspectNet 2022 ResNet-50 Mono 96.16 93.87 88.17 87.17 76.57 68.42 86.15 77.11 71.24
@@
[51] Gaussian YOLOv3 2019 Darknet-53 Mono 90.61 90.20 81.19 87.84 79.57 72.30 89.31 81.30 80.20
@@
[51] Gaussian YOLOv3 2019 Darknet-53 Mono 98.74 90.48 89.47 87.85 79.96 76.81 90.08 86.59 81.09
@@
[289] ASSD 2019 ResNet-101 Mono 89.28 89.95 82.11 69.07 62.49 60.18 75.23 76.16 72.83
@@
[307] RefineDet 2018 VGG-16 Mono 98.96 90.44 88.82 84.40 77.44 73.52 86.33 80.22 79.15
@@
[250] SqueezeDet 2017 SqueezeNet Mono 90.20 84.70 73.90 77.10 68.30 65.80 82.90 75.40 72.10
@@
From Table 14, we provide an in-depth discussion of two selected models: AspectNet [140] and SINet [91], which achieved the highest performance among all evaluated methods, and Gaussian YOLOv3 [51], one of the most widely used models in the context of autonomous driving. AspectNet [140] approaches object detection tasks uniquely with an anchor-free detection scheme. It sets itself apart from comparable models by adding a special head that predicts object aspect ratios. Rather than tackline classification and regression features separately, the network combines them before reaching the aspect module, which allows geometric details to play a role in both areas. This additional branch is designed with just a few convolutional layers, ensuring it does not slow inference time too much. Another notable change is how positive and negative training samples are defined: the criteria for matching are tweaked to tackle the class imbalance that can hurt performance in single-stage models. These architectural decisions enhance the detector’s ability to recognize elongated or skewed objects, as the receptive field is shaped better to fit the actual geometry of the detected objects.
@@
Gaussian YOLOv3 [51] keeps the overall layout of YOLOv3 intact, with three detection scales, feature aggregation from the backbone, and dense predictions. However, it changes the way bounding boxes are parameterized. Each coordinate is modeled as a Gaussian variable with its variance, meaning the network outputs not only position estimates but also an uncertainty term for each. These variances are fed into a modified regression loss based on negative log-likelihood, and they also influence the suppression of overlapping boxes at inference. The result is a system that can flag low-confidence localizations without discarding high-confidence detections that fall near ambiguous boundaries. Compared with AspectNet’s emphasis on encoding geometric priors into a deterministic pipeline, Gaussian YOLOv3 takes a probabilistic route to preserve the speed of a one-stage detector while offering a measure of confidence that can be applied to downstream tasks.
@@
Table 15 focuses on 3D object detection methods that use camera-based KITTI data, but generate 3D bounding boxes from 2D image inputs. From Table 15, we provide a detailed discussion of four selected models: M3D-RPN [22], recognized as one of the most widely adopted models in this field; MonoDGP [182], which is the latest released method; MonoTAKD [143], which achieved the highest results in $AP_{BEV}$ CarCar and $Pedestrian$ evaluations; and MonoDiff [189] which achieved the highest results in $AP_{3D}$ CarCar evaluation on the KITTI dataset. It should be mentioned that the values highlighted in red represent the best performance achieved for each evaluation metric across all listed methods.
@@
The Monocular 3D Region Proposal Network (M3D-RPN) [22] presents an integrated approach for generating 2D and 3D object proposals from a single RGB image. Built upon a DenseNet-121 backbone modified with dilated convolutions to preserve a stride of 16, the network branches into two complementary pathways: one employing standard spatially invariant convolutions to capture global context, and another utilizing depth-aware convolutions to encode row-dependent geometric cues. In the latter, feature maps are divided into horizontal bins, with each bin processed by distinct convolutional kernels to reflect perspective variations inherent to urban driving scenes. Anchors are designed to carry 2D and 3D parameters, initialized from dataset-derived statistics to provide strong geometric priors. Predictions from the two branches are combined to regress full 3D bounding box parameters, followed by a lightweight post-optimization step that refines object orientation through projection consistency between 3D boxes and their 2D counterparts. M3D-RPN reduces complexity while maintaining accuracy by consolidating proposal generation into a single end-to-end stage.
@@
MonoDGP [182] employs a transformer-based architecture that unites geometric priors with a novel perspective-invariant error formulation for depth refinement. A ResNet-50 backbone extracts multi-scale features, which are enhanced by a Region Segmentation Head (RSH) to emphasize foreground content and embed segment labels distinguishing object regions from background. This network’s detection is decoupled into two parallel decoding processes: a 2D visual decoder for initializing queries and reference points from appearance cues and a 3D depth-guided decoder that integrates visual and depth features to model spatial relationships. Rather than relying on direct depth prediction or multi-branch fusion, MonoDGP computes geometric depth using the projection formula and introduces an explicit geometry error term to correct systematic offsets caused by perspective and object surface effects. With its stable and compact statistical distribution, this error term reduces training complexity while improving depth accuracy. Combining query decoupling allows the method to achieve improved convergence stability and precise 3D localization without requiring auxiliary data sources.